OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -1 + 2 + 5 + 8 + ... + 44 PLEASE HELP, THANKS!

OpenStudy (kc_kennylau):

Find the formula for the sequence first

OpenStudy (anonymous):

Okay, let me work on that.

OpenStudy (anonymous):

I am at a loss right now, I really do not understand how to go about this problem. Could you walk me through it, if you are willing?

OpenStudy (kc_kennylau):

Find the common difference first

OpenStudy (anonymous):

d=3

OpenStudy (kc_kennylau):

Find the difference between the common difference and the first term

OpenStudy (anonymous):

as in 3-(-1)= 4?

OpenStudy (kc_kennylau):

Yep

OpenStudy (kc_kennylau):

Therefore \(\Large a_n=-3n+4\)

OpenStudy (kc_kennylau):

What does n equal when: (1) \(a_n=-1\)? (2) \(a_n=44\)?

OpenStudy (anonymous):

(1) n= 7 (2) n= -128 ?

OpenStudy (kc_kennylau):

You reversed them, I'm asking the opposite

OpenStudy (anonymous):

I apologize if I am wrong; I am sincerely confused when it comes to these problems. I will try again.

OpenStudy (kc_kennylau):

You thought I gave you values of n and you found a_n, where in fact I gave you values of a_n and you find n

OpenStudy (anonymous):

So would I set up equations as follows? : (1) -1=-3n+4 (2) 44=-3n+4

OpenStudy (kc_kennylau):

Yep

OpenStudy (kc_kennylau):

Because you have the formula of the sequence

OpenStudy (kc_kennylau):

You want to know where it starts and where it ends

OpenStudy (kc_kennylau):

-1 is the start and 44 is the end

OpenStudy (anonymous):

Oh, I see my mistake. So then we have the values: (1) 5/3 (2) -(40/3) I do understand the start and ending points though.

OpenStudy (anonymous):

Or do I actually not solve out the equations and I am making a fool of myself?

OpenStudy (kc_kennylau):

Wait my bad it should be \(\Large a_n=3n+4\) and yes you do solve the equation

OpenStudy (kc_kennylau):

You want to know which n corresponds to -1 (the first term) and which n corresponds to 44 (the last term)

OpenStudy (kc_kennylau):

And you aren't making a fool of yourself

OpenStudy (anonymous):

(Thank you holding in here with me; it's greatly appreciated.) (1) -(5/3) and this would correspond to the first term, correct? (2) 40/3 and this would correspond to the second term?

OpenStudy (kc_kennylau):

Please solve the equations again

OpenStudy (kc_kennylau):

Because of my bad I modified the equationd

OpenStudy (kc_kennylau):

I'll explain the significance later

OpenStudy (anonymous):

(1) -1= 3n+4 (2) 44= 3n+4 Wouldn't their answers be the same as before, but just have opposite signs?

OpenStudy (kc_kennylau):

Sorry my bad again... It should be \(a_n=3n-4\)

OpenStudy (anonymous):

Oh, and it's okay. (1) -1= 3n-4 = 1 (2) 44= 3n-4 = 16

OpenStudy (kc_kennylau):

Therefore when n=1, a_n=-1 When n=16, a_n=44

OpenStudy (kc_kennylau):

Therefore the 1st term is -1 and the 16th term is 44

OpenStudy (kc_kennylau):

Therefore it's \(\displaystyle\Large\sum_{n=1}^{16}(3n-4)\)

OpenStudy (anonymous):

This makes perfect sense! I understand the process of writing out eh summation notation, thank you! There's just one problem; these are the only answer choices: \[\sum_{n=0}^{15}-3n\] \[\sum_{n=0}^{15}(-1+3n)\] \[\sum_{n=0}^{\infty}-3n\] \[\sum_{n=0}^{\infty}(-1+3n)\]

OpenStudy (kc_kennylau):

Oh that's just the same

OpenStudy (kc_kennylau):

\[\sum_{n=1}^{16}(3n-4)=\sum_{n=0}^{15}(3n-1)\]

OpenStudy (anonymous):

Oh, I was worried there for a bit. Thank you so much for your help; you have been of great service to me! Thanks a ton!! 

OpenStudy (kc_kennylau):

No problem